On March 1, 2018, you start “Road to Broad T-Shirts” in a kiosk outside the Business College. The following events occurred during the month of March:

Investment: You use your savings of $60,000 to start the business as a corporation as the only stockholder.

Kiosk Purchase: You buy a kiosk for $12,600 cash. The kiosk will be useful for six months and will have no value at the end of the sixth month.

Labor Cost: You hire your friend, Joey, to work the kiosk. You pay him $800 on March 31, for his work during the month of March.

Cell Phone Bill: Your first cell phone bill arrived at the end of March totaling $125; which you will pay on the first day of April.

T-Shirts: On March 1 you buy 900 T-shirts, paid in cash on delivery, each for $9. During March, you sell 600 of the T-shirts, each for $21, selling 450 for cash and 150 to a student group on credit. The student group plans to pay you back by April 15.

a). At the end of March, how much cash did Road to Broad T-Shirts have?

b). How well did Road to Broad T-Shirts perform in March? In other words, how profitable was it?

c). What were Road to Broad T-Shirts' resources and obligations at the end of March?

d). You should find the net resources and obligations at the end of March is $64,175. How do you explain the source of this amount?
Answer #1
Answer to part A Cash Balance of Road to Broad T-Shirts
Cash Account
To Cash Broughtin to business 60000
To Sales 9450 By Kiosk Purchase 12600
By wage Paid 800
By Purchses(suppliers) 8100
By Balance 47950
69450 69450

So Cash Balance for end March is $47950

Answer to B part-Profitability of Road to braod T _shirts
Road to Broad T-Shirts
PROFIT & LOSS ACCOUNT FOR THE MARCH
PARTICULARS $ PARTICULARS $
To Purchases        8,100 By sales       12,600
To Wages           800 Byclosing stock        2,700
To Telephone Bill Payable           125
To Depreciation on Kiosk        2,100
To Net Profit        4,175
      15,300       15,300

So the profit for march is $4175

Answer to Part c- Resource and obligations

Road To Broad T-shirts
BALANCE SHEET( RESOURCE AND OBLIGATIONS) AS ON 31st march
LIABILITIES/OBLIGATIONS $ ASSETS/RESOURCE $
Capital Account Kiosk 10500
investment 60000 Cash & Bank Balance 58,450
Profit 4175 64,175 Closing Stock    2,700
Telephone Bill payable         125 Sundry Receivables    3,150
TOTAL 64,300 TOTAL 64,300
Workings
Sundry Receivables 150*21=3150
Kiosk 12600-2100=10500
closing stock 300*9=2700

Part D Net resource = Total resource- Obligations
=64300-125  

= $ 64175

This $ 64175 is as a result of Intial investment 60000 and earnings From Business 4175